OpenStudy (anonymous):

Rewrite 2y=3x+12 in standard form

OpenStudy (anonymous):

show me an equation for standard form.

OpenStudy (austinl):

Ax+By=C C usually =1

OpenStudy (anonymous):

then rearrange 2y-3x=12 -3x+2y=12

OpenStudy (anonymous):

well i just cant do standad form wid this type of equation that y i need help

OpenStudy (anonymous):

if you want 12 to equal 1 divide both sides by 12 -3x+2y=12 (-3x+2y)/12 =1

OpenStudy (anonymous):

you means standard form or y=mx+b form?

OpenStudy (anonymous):

(-3x+2y)/12 =1 -3/12x+2/12y=1

OpenStudy (anonymous):

stantdard in general because im very good at doing y = mx + b but im so confused wid standardforms

OpenStudy (anonymous):

3x - 2y = 12 right???

OpenStudy (anonymous):

Ax+By=C C usually =1 2y=3x+12 2y-3x=12 subtract 3x both sides -3x+2y=12 rearanged left hand side (-3x+2y)/12=1 divide both sides by 12 -3/12x+2/12y=1